OpenStudy (ashleyy_smith):

I'll give a medal to whoever answers this, Two cars leave towns 800 kilometers apart at the same time and travel toward each other. One car's rate is 16 kilometers per hour more than the other's. If they meet in 4 hours, what is the rate of the faster car?

OpenStudy (tylermckinney16):

Idk lol. but i have a friend whos good at these so ill show him lol and get back to ya if i figure it out lol.

OpenStudy (ashleyy_smith):

Oh alright :o

OpenStudy (ashleyy_smith):

@vishweshshrimali5

OpenStudy (tylermckinney16):

Yeah lol sense we can give medals ill also give ya a medal idk why just will lol ill tag some of my friends that are on here in this as well lol. there really good.

OpenStudy (vishweshshrimali5):

We will continue from where we left earlier

OpenStudy (ashleyy_smith):

alright

OpenStudy (tylermckinney16):

@Directrix @Astrophysics @ILovePuppiesLol

OpenStudy (vishweshshrimali5):

Give me a sec...

OpenStudy (ashleyy_smith):

alright (:

OpenStudy (tylermckinney16):

Well i tagged some friends who might be able to help :) u have a great day :)

OpenStudy (ashleyy_smith):

you too (:

OpenStudy (tylermckinney16):

Thanks :)

OpenStudy (vishweshshrimali5):

v2 = v1 + 16... So, the equation should be: d1 + d2 = 800 => v2*t + v1 * t = 800 => (v1 + 16) * t + v1 * t = 800 We know t = 4 Thus: (v1 + 16)*4 + v1*4 = 800 => v1 * 4 + 16 * 4 + v1 * 4 = 800 => 8 * v1 + 16 * 4 = 800

OpenStudy (vishweshshrimali5):

We were here I guess?

OpenStudy (ashleyy_smith):

yes

OpenStudy (vishweshshrimali5):

Great so only thing left is to solve the equations right?

OpenStudy (ashleyy_smith):

yup

OpenStudy (vishweshshrimali5):

Okays! So we will start with this equation: 8 * v1 + 16 * 4 = 800 Now tell me what is 16*4 ?

OpenStudy (ashleyy_smith):

64

OpenStudy (vishweshshrimali5):

Great so our equation becomes: 8*v1 + 64 = 800 => 8 * v1 = 800 - 64 Now, what is 800 - 64?

OpenStudy (ashleyy_smith):

736

OpenStudy (vishweshshrimali5):

Very nice... so our equation now becomes... 8 * v1 = 736 Now what is the value of v1 from here ?

OpenStudy (ashleyy_smith):

92?

OpenStudy (vishweshshrimali5):

Good :)

OpenStudy (vishweshshrimali5):

Now we use our final equation v2 = v1 + 16 => v2 = 92 + 16 So what do you get ?

OpenStudy (tylermckinney16):

Man idk how u do that all that math kills my head lol.

OpenStudy (ashleyy_smith):

108

OpenStudy (vishweshshrimali5):

Great! So here you go... the answer is 108 km/hr

OpenStudy (vishweshshrimali5):

Any doubt ?

OpenStudy (ashleyy_smith):

(: nope I'm done lol thanks to u

OpenStudy (vishweshshrimali5):

:) Glad to be of help